Let f and g be functions from the positive real numbers to positive real numbers defined by f(x) = [2x]

g(x) = x2.2 Calculate f o g and g o f.


1
Expert's answer
2021-09-20T15:22:19-0400

The integral part or integer part of "x," often denoted "[x]" is usually defined as the "\lfloor x\rfloor" if "x" is nonnegative, and "\lceil x\rceil" otherwise.



"f\circ g=[g]=[2x^2]=\lfloor 2x^2\rfloor, since \ 2x^2\geq0,x\in\R"

"g\circ f=(f)^2=([2x])^2"